OpenStudy (anonymous):

If 3x^2-2x+7=0, then (x-1/3)^2= A 20/9 B 7/9 C -7/9 D -8/9 E -20/9

OpenStudy (triciaal):

one approach expand (x-1/3)^2 and substitute 3x^2 -2x+ 7 = 0 then simplify

OpenStudy (anonymous):

Ok so should i substitute (x-1/3)^2 into the other one?

OpenStudy (anonymous):

is this what the first one looks like expanded? x^2-2/3+1/9

OpenStudy (triciaal):

another approach to factor complete the square 3x^2 -2x + 7 = 0 divide by 3 (x-b/a)^2 = (x-2/6)^2 = (x-1/3)^2 you had to add 1/9

OpenStudy (triciaal):

*should be b/2a

OpenStudy (anonymous):

To factor that equation you said i should divide by three first?

OpenStudy (triciaal):

you will get the same result 3x^2-2x + 7 = 0 x^2 -2/3 + 7/3 =0 (x-2/6)^2 + 1/9 + 7/3 = 0 (x-1/3)^2 = 0-1/9-7/3 = do you see E

OpenStudy (anonymous):

Ohh yes i see that thank you so much you're awesome :)

OpenStudy (triciaal):

wait i think i have a calculation error

OpenStudy (anonymous):

i have a question did u substitute the other equation or where did u get 2/6?

OpenStudy (triciaal):

(x-2/6)^2 + 1/9 + 7/3 = 0 I added 1/9 to the left so i need 1/9 on the right

OpenStudy (anonymous):

Oh thats right

OpenStudy (anonymous):

Alright i'll do this once again and see if i can get the same answer :)

OpenStudy (triciaal):

it's ok I actually did it but i did not write it on the entry posted 1/9 -7/3 = 1/9-21/9 = -20/9 answer E

OpenStudy (triciaal):

when you complete the square x-b/2a a = 3 b = 2 2*a = 2*3 = 6

OpenStudy (triciaal):

(a +b)^2 = a^2 + 2ab + b^2 you have 2ab when expanded

OpenStudy (anonymous):

thanks a lot you're a great help :)